OpenStudy (anonymous):

Factor completely. x^3+5x^2+6x

OpenStudy (anonymous):

give me sum to drink ha

OpenStudy (anonymous):

first take out x = x(x^2 + 5x +6) = x(x+3)(x+2)

OpenStudy (anonymous):

heres a coke delaine!

OpenStudy (anonymous):

\[given: x^3 + 5x^2 + 6x x(x^2 + 5x + 6)\] \[The \ factors \ are\ x = 0,\ or\ x= -2,\ or \ x= -3\]

OpenStudy (anonymous):

\[medal\ plz\]